The larger of two numbers is twice the smaller. I the larger is decreased by ten, the result is five more than the smaller. Find the numbers.
Answer by JimboP1977(308)

x = larger number
y = smaller number
x = 2y (1)
x-10 = 5y (2)
(2) - (1)
-10 = 3y
y = -10/3
x = -20/3